# Gravitational-Wave Signatures of Highly Eccentric Stellar-Mass Binary Black Holes in Galactic Nuclei

### v1

@@ Abstract
A significant fraction of compact-object mergers in galactic nuclei are expected to be eccentric in the Laser-Interferometer-Space-Antenna (LISA) frequency sensitivity range, $10^{-4}-10^{-1}\ \rm Hz$. Several compact binaries detected by the LIGO-Virgo-KAGRA Collaboration may retain hints of residual eccentricity at $10$ Hz, suggesting dynamical or triple origins for a significant fraction of the gravitational-wave-observable population. In triple systems, von-Zeipel-Lidov-Kozai oscillations perturb both the eccentricity and the argument of pericentre, $\omega$, of the inner black hole binary. The latter could be fully circulating, where $\omega$ cycles through $2\pi$, or may librate, with $\omega$ ranges about a fixed value with small or large variation. We use TSUNAMI, a regularised N-body code with up to 3.5 post-Newtonian (PN) term corrections, to identify four different families of orbits: (i) circulating, (ii) small and (iii) large amplitude librating, and (iv) merging orbits. We develop and demonstrate a new method to construct gravitational waveforms using the quadrupole formula utilising the instantaneous total acceleration of each binary component in TSUNAMI. We show that the four orbital families have distinct waveform phenomenologies, enabling them to be distinguished if observed in LISA. The orbits are also distinguishable from an isolated binary or from a binary perturbed by a different tertiary orbit, even if the secular timescale is the same. Future burst timing models will be able to distinguish the different orbital configurations. For efficient binary formation, about $\sim 50$ binaries can have librating orbits in the Galactic Centre.
@@ 1 Introduction
In contrast, dynamical formation occurs through gravitational interactions in densely-populated stellar environments. These interactions can assemble binaries with eccentric orbits, large or misaligned spins, and component masses exceeding the pair-instability limit through hierarchical mergers (see Mandel & Broekgaarden, 2022, for details). Dynamical channels include a range of environments, such as open clusters (Kumamoto et al., 2019; Trani et al., 2021), globular clusters (Rodriguez et al., 2016; Samsing, 2018), and nuclear star clusters (Antonini & Perets, 2012; Grishin et al., 2018; Hoang et al., 2018). Black holes embedded in the dense accretion discs of active galactic nuclei (AGN) may also merge at significant rates (Stone et al., 2017; Tagawa et al., 2020; Samsing et al., 2022; Trani et al., 2024), although the relevant physical processes remain uncertain (Gilbaum & Stone, 2022; Grishin et al., 2024; Gilbaum et al., 2025; Trani & Di Cintio, 2025).
@@ 2.1.2 Inner binary orbits
Consider a binary of masses $m_{1}$ and $m_{2}$ (or total mass $m_{\rm bin}=m_{1}+m_{2}$) with semi-major axis $a_{1}$ and initial eccentricity $e_{1}$. The centre of mass revolves around a central SMBH of mass $m_{3}\gg m_{1}+m_{2}$ on semi-major axis $a_{2}$ and eccentricity $e_{2}$. The mutual inclination between the two orbital planes is $\iota_{\rm mut}$. Fig. 1 sketches the orbital elements of the hierarchical triple system. If the tertiary is too close, the binary is dynamically unstable. Many studies have been devoted to stability of triple systems (e.g., Mardling & Aarseth, 2001; Grishin et al., 2017; He & Petrovich, 2018; Tory et al., 2022; Lalande & Trani, 2022; Vynatheya et al., 2022; Hayashi et al., 2022, 2023). In our case, the measure of stability is encapsulated in the Hill radius,
@@
To summarise, the range of inner binary orbits where the SMBH is able to perturb the orbit but still keep it stable spans around one order of magnitude, $0.04\ \rm au\lesssim a_{1}\rm\ \lesssim 0.4\ \rm au$. The upper limit is due to the Hill stability $a_{1}\lesssim 0.5r_{\rm H}$ (Eq. 1), while the lower limit is due to transition to 1PN-GR dominated precession (Eq. 2), where the binary will effectively evolve in in isolation.
@@ 2.2 Properties of highly inclined triple systems
For a dynamically stable binary dominated by the SMBH over GR ($\epsilon_{\rm GR}\leq 1)$, quasi-periodic ZLK oscillations occur if the mutual inclination $\iota_{\rm mut}$ is large enough. The almost co-planar orbits, $\omega$ is always circulating and completes a full revolution in a secular timescale. Beyond a critical inclination, a bifurcation in the phase portrait leads to a fixed point in $e_{1}$-$\omega$ space and orbits can either circulate (where $\omega$ completes a full revolution) or librate (where $\omega$ is limited in range, see e.g., Naoz, 2016). This is analogous to the pendulum model where $\omega$ plays the role of the angle and $e_{1}$ is related to the ’momentum’ (see Basha et al., 2025, for a detailed analysis). There is a critical curve that lies on the unstable fixed point (where the pendulum is balanced on the top) that separates between circulating and librating orbits, which is called the separatrix.
ZLK dynamics do not have fixed points in $e_{1}$ - $\omega$ phase space, and $\omega$ always complete a full circulation. Beyond a critical inclination, a bifurcation in the phase portrait leads to a fixed point and orbits can either circulate (where $\omega$ completes a full revolution) or librate (where $\omega$ is limited in range, see e.g., Naoz, 2016). Hansen & Naoz (2020) studied the structure of fixed and saddle points of the hierarchical three body problem to octupole order where the additional dependence on the outer eccentricity and the nodal angle lead to a diverse orrery of families of fixed and saddle points. Here we focus on circular outer orbit with equal mass inner binaries, thus the octupole term is absent, but we take into account the corrections due to mild hierarchies, as explained below:
@@
Mild hierarchies where the inner binary spans a significant fraction of the Hill sphere have extra corrections proportional to the period ratio $\epsilon_{\rm SA}\sim P_{1}/P_{2}$, in the form of the Brown Hamiltonian (Tremaine, 2023). The main effect is shifting the locations of the bifurcation points and the extent of the maximal eccentricity, but adding the Brown Hamiltonian does not change the morphology of the phase space. The analytical expression for the fixed points are given by Grishin (2024a):
@@ 2.3 Numerical integrations and methods
In all the following integrations, we use the post-Newtonian N-body code TSUNAMI (Trani & Spera, 2023) to integrate a three-body system. TSUNAMI is fast, implements chain-regularisation, and has PN terms up to 3.5PN (Bulirsch & Stoer, 1966; Mikkola & Aarseth, 1993; Mikkola & Tanikawa, 1999).
@@
where the time-derivatives are the instantaneous velocities and accelerations from TSUNAMI  [^1] In principle, additional acceleration can rise from interactions with the gas in AGN discs (Samsing et al., 2022; Gangardt et al., 2024) or during inspiral in a stellar envelope (Ginat et al., 2020; Trani et al., 2022a) where the torque prescription is fitted to detailed hydrodynamical simulations..
@@
Table 1: Initial conditions and final integration times for the four orbital families. The BH masses are always qual to $m_{1}=m_{2}=20M_{\odot}$. The inner semi-major axis is always $a_{1}=0.15\ \rm au$.
@@ 3 Orbital families
Before going to the detailed dynamics of each orbital family, we provide a brief overview and the key characteristic of each family.
In Figure 2, we plot the correlation between $\omega$ and the GW frequency $f_{\rm GW}$ (Eq. 5) for the different families. Small libation occurs when the orbit is initialised with initial conditions at the fixed point (the small libration due to higher order effects, blue dots). For initial conditions further from the fixed point, but still ones that allow libration in $\omega$, the libration amplitude of $\omega$ and consequently of $e_{1}$ and $f_{\rm GW}$ is larger (black triangles). For an orbit outside the separatrix, $\omega$ completes a full circulation. The GW frequency becomes LISA-observable around $0.01$ Hz for some limited values or $\omega$ (red dots). For an initially librating orbits with sufficiently large eccentricity, GW dissipation at the pericentre significantly shifts the curves to a new energy curve. Eventually the orbit is tight enough and begins to decouple from the tertiary. At some point $\omega$ starts to circulate and eventually is dominated by GR 1PN precession without changing the eccentricity (green triangles). This orbit regularly achieves GW frequency above $0.1\ \rm Hz$.
@@
Figure 4: Representative of the large amplitude libration orbital family. Similar to Fig. 3, bit with different initial conditions (see table 1).
@@
Figure 6: Representative of the librating and merging family of orbits. The panels are similar to Fig. 3. The initial conditions are described in table 1.
@@ 3.1 Circulating orbits
The only example of a population of highly inclined hierarchical triples with extreme mass ratios is the population of irregular satellites around giant planets in the Solar System. Out of the $228$ irregular satellites [^2] By irregular we mean satellites with orbit beyond the Laplace radius (Tremaine, 2023), which are believed to be formed by capture., only $13$ are found in ZLK libration, the rest are circulating (Grishin, 2024b). Theoretically, the area in $e_{1}$-$\omega$ space space is larger for circulating orbit, with the actual ratio depending on $j_{z}$. Thus, the circulating family of orbits is the most common. Fig. 3 depicts the orbital evolution of a representative orbit. The orbit starts from a relatively low/moderate eccentricity (in this case $e_{1}=0.5$, top left) and grows to large values. The mutual inclination is sharply decreasing in a correlated manner that preserves $j_{z}$ (on average; middle left). The argument of pericentre $\omega$ is circulating and completes a full revolution (bottom left).
The top middle panel shows the instantaneous separation ${\rm d}r$ and the pericentre $a_{1}(1-e_{1})$ which also changes by almost three orders of magnitude. The central panels show the evolution of the peak frequency $f_{\rm GW}$ with time. The peak frequency grows by almost three orders of magnitude for the circulating family. The middle bottom panel show both strain polarisations $h_{+}$ and $h_{\times}$ for $D=8\ \rm kpc$, $\theta,\phi=0$ (head-on GW propagation). We see that thee strain is highly bursty and peaks to strain magnitudes of about $10^{-19}$ for a short duration during the highly eccentric phase.
@@ 3.2 Large amplitude librating orbits
Fig. 4 shows an example of an orbit that librates with a large amplitude $\omega$ between $\sim 0.75$ and $2.25$ rad, a range of almost $\pi/2$ radians (or 90 degrees) (top right). The only difference between this orbit and the circulating orbit is a different starting point of $\omega$. The peak frequency $f_{\rm GW}$ also changes by over two orders of magnitude, and the maximum peak is again achieved at $\omega\approx\pi/2$. The phase portrait of $e_{1}$-$\omega$ is now on a closed loop, but the behaviour is similar to the circulating orbit where $\omega$ sweeps quickly in the $e_{1}\sim\omega$ phase space. This is associated with the greatest loss of GW radiation, which is also indicated in the peak of the strain polarisations. Overall the large libration orbital family is similar to the circulating one. This is because these orbits are relatively close to the separatrix, which is the critical curve that separates circulating and librating orbits and crosses the unstable (hyperbolic) fixed point of $e_{1}=0,\omega=0$, similar to the pendulum. (see details e.g., in Grishin, 2024a).
@@ 3.3 Small amplitude librating orbits
Fig. 5 shows the orbital evolution of a small amplitude librating orbit. We see that the orbital elements vary little, and the variation in $f_{\rm GW}$ varies by $0.4$ dex. The argument of pericentre is confined to a small range of $\sim 7.5$ degrees. The system stays at relatively high eccentricity, and the waveform has relatively regular bursts on the inner period timescale $P_{1}\approx 0.0092\ \rm yr$, which is different from the other systems that experience quiescent phases over the secular timescale of $\sim 1-2\ \rm yr$. The phase portrait shows clear motion of small librations around a fixed point. We will examine the small amplitude libration orbit, the different timescale involved (inner orbit $P_{1}\ll$ outer orbit $P_{2}\ll$ secular time $t_{\rm sec}$ and the comparison to an isolated binary evolution in sec. 4.3.

### v2

@@ Abstract
A significant fraction of compact-object mergers in galactic nuclei are expected to be eccentric in the Laser Interferometer Space Antenna (LISA) frequency sensitivity range, $10^{-4}-10^{-1}\ \rm Hz$. Several compact binaries detected by the LIGO-Virgo-KAGRA Collaboration may retain hints of residual eccentricity at $\sim 10$ Hz, suggesting dynamical or triple origins for a significant fraction of the gravitational-wave-observable population. In triple systems, von-Zeipel-Lidov-Kozai oscillations perturb both the eccentricity and the argument of pericentre, $\omega$, of the inner black hole binary. The latter could be fully circulating, where $\omega$ cycles through $2\pi$, or may librate, with $\omega$ ranges about a fixed value with small or large variation. We use TSUNAMI, a regularised N-body code with up to 3.5 post-Newtonian (PN) term corrections, to identify four different families of orbits: (i) circulating, (ii) small and (iii) large amplitude librating, and (iv) merging orbits. We develop and demonstrate a new method to construct gravitational waveforms using the quadrupole formula utilising the instantaneous total acceleration of each binary component in TSUNAMI. We show that the four orbital families have distinct waveform phenomenologies, enabling them to be distinguished if observed in LISA. The orbits are also distinguishable from an isolated binary or from a binary perturbed by a different tertiary orbit, even if the secular timescale is the same. Future burst timing models will be able to distinguish the different orbital configurations. For efficient binary formation, about $\sim 1000$ binaries can have highly eccentric, librating orbits in the Galactic Centre.
@@ 1 Introduction
In contrast, dynamical formation occurs through gravitational interactions in densely-populated stellar environments. These interactions can assemble binaries with eccentric orbits, large or misaligned spins, and component masses exceeding the pair-instability limit through hierarchical mergers (see Mandel & Broekgaarden, 2022, for details). Dynamical channels include a range of environments, such as open clusters (Kumamoto et al., 2019; Trani et al., 2021), globular clusters (Rodriguez et al., 2016; Samsing, 2018), and nuclear star clusters (Antonini & Perets, 2012; Grishin et al., 2018; Hoang et al., 2018). Black holes embedded in the dense accretion discs of active galactic nuclei (AGN) may also merge at significant rates (Stone et al., 2017; Tagawa et al., 2020; Samsing et al., 2022; Trani et al., 2024), although the relevant physical processes remain uncertain (Gilbaum & Stone, 2022; Grishin et al., 2024; Gilbaum et al., 2025; Moncrieff et al., 2026).
@@ 2.1.2 Inner binary orbits
Consider a binary of masses $m_{1}$ and $m_{2}$ (or total mass $m_{\rm bin}=m_{1}+m_{2}$) with semi-major axis $a_{1}$ and initial eccentricity $e_{1}$. The centre of mass revolves around a central SMBH of mass $m_{3}\gg m_{1}+m_{2}$ on semi-major axis $a_{2}$ and eccentricity $e_{2}$. The mutual inclination between the two orbital planes is $\iota_{\rm mut}$. Figure 1 sketches the orbital elements of the hierarchical triple system. If the tertiary is too close, the binary is dynamically unstable. Many studies have been devoted to stability of triple systems (e.g., Mardling & Aarseth, 2001; Grishin et al., 2017; He & Petrovich, 2018; Tory et al., 2022; Lalande & Trani, 2022; Vynatheya et al., 2022; Hayashi et al., 2022, 2023). In our case, the measure of stability is encapsulated in the Hill radius,
@@
To summarise, the range of inner binary orbits where the SMBH is able to perturb the orbit but still keep it stable spans around one order of magnitude, $0.04\ \rm au\lesssim a_{1}\rm\ \lesssim 0.4\ \rm au$. The upper limit is due to the Hill stability $a_{1}\lesssim 0.5r_{\rm H}$ (Eq. 1), while the lower limit is due to transition to 1PN-GR dominated precession (Eq. 2), where the binary will effectively evolve in isolation.
@@ 2.2 Properties of highly inclined triple systems
For a dynamically stable binary dominated by the SMBH over GR ($\epsilon_{\rm GR}\leq 1)$, quasi-periodic ZLK oscillations occur if the mutual inclination $\iota_{\rm mut}$ is large enough. For almost co-planar orbits, $\omega$ is always circulating and completes a full revolution in a secular timescale. Beyond a critical inclination, a bifurcation in the phase portrait leads to a fixed point in $e_{1}$-$\omega$ space and orbits can either circulate (where $\omega$ completes a full revolution) or librate (where $\omega$ is limited in range, see e.g., Naoz, 2016). This is somewhat analogous to the pendulum model (see Basha et al., 2025, for a detailed analysis) [^1] The ’angle’ in Basha et al. (2025) is not $\omega$, but defined via a combination of the eccentricity and angular momentum vectors (see their Eq. 25). The ’momentum’ is proportional to the $z$ component of the eccentricity vector, while the potential energy (i.e. the ’height’ of the pendulum) is related to the eccentricity $e_{1}$. Moreover, circulation of $\omega$ corresponds to a libration in the pendulum model, and vice versa (see their Figure 1 and 2).. There is a critical curve that lies on the unstable fixed point (where the pendulum is balanced on the top) that separates between circulating and librating orbits, which is called the separatrix. Hansen & Naoz (2020) studied the structure of fixed and saddle points of the hierarchical three body problem to octupole order where the additional dependence on the outer eccentricity and the nodal angle lead to a diverse orrery of families of fixed and saddle points. Here we focus on circular outer orbit with equal mass inner binaries, thus the octupole term is absent, but we take into account the corrections due to mild hierarchies, as explained below:
@@
Mild hierarchies where the inner binary spans a significant fraction of the Hill sphere have extra corrections proportional to the period ratio $\epsilon_{\rm SA}$ in the form of the Brown Hamiltonian (Tremaine, 2023)
$\epsilon_{\rm SA}=\left(\frac{a_{1}}{a_{2}(1-e_{2}^{2})}\right)^{3/2}\left(\frac{M_{\rm SMBH}}{m_{0}+m_{1}}\right)^{1/2}\approx\frac{P_{1}}{P_{2}}.$ (3)
The main effect is shifting the locations of the bifurcation points and the extent of the maximal eccentricity, but adding the Brown Hamiltonian does not change the morphology of the phase space. The analytical expression for the fixed points are given by Grishin (2024a):
@@
For the highly inclined orbits we consider here, the attained eccentricities are high, and the secular approximation breaks down (Antonini & Perets, 2012; Grishin et al., 2017, 2018). Fragione et al. (2019) shown that direct N-body integrations produce 5105-10 times more mergers (and many eccentric ones) over the secular approach of Hoang et al. (2018) for the same initial conditions. Mangipudi et al. (2022) have studied the modified expression for the maximal eccentricities and compared the results with TSUNAMI and found that for the types of orbits considered here where $\epsilon_{\rm SA}\approx 0.05$, the secular approach would fail to capture the highly eccentric bursts. Thus, the direct approach of computing the strain from the accelerations greatly improves upon existing tools of inferring the strain from the secular theory alone (e.g., Knee et al., 2024).
@@ 2.3 Numerical integrations and methods
In all the following integrations, we use the post-Newtonian $N$-body code TSUNAMI (Trani & Spera, 2023) to evolve a three-body system. TSUNAMI combines Bulirsch–Stoer extrapolation (Bulirsch & Stoer, 1966), a regularized leapfrog algorithm (Mikkola & Tanikawa, 1999), and chain coordinates (Mikkola & Aarseth, 1993) to achieve high accuracy and computational efficiency. The extrapolation algorithms uses an adaptive timestep based on convergence rejection with a relative tolerance of $10^{-13}$ (Press, Flannery, & Teukolsky, 1986). The code implements post-Newtonian acceleration corrections up to order 3.5 in $v/c$, formulated in the modified harmonic coordinate system (Blanchet, 2024). Gravitational waveforms are computed from the positions, velocities, and accelerations using the quadrupole formula (e.g. Maggiore, 2007), where the accelerations consistently account for the total forces acting on the binary components.
@@
where the time-derivatives are the instantaneous velocities and accelerations between particles as calculated as in TSUNAMI  [^2] In principle, this term may also include additional accelerations arising from interactions with gas, such as in AGN discs (Samsing et al., 2022; Gangardt et al., 2024) or during inspiral within a stellar envelope (Ginat et al., 2020; Trani et al., 2022a), where torque prescriptions are calibrated against hydrodynamical simulations. However, such effects are not relevant for the systems considered in this work..
@@
Table 1: Initial conditions and final integration times for the four orbital families. The BH masses are always equal to $m_{1}=m_{2}=20M_{\odot}$. The inner semi-major axis is always $a_{1}=0.15\ \rm au$.
@@ 3 Orbital families
Before going to the detailed dynamics of each orbital family, we provide a brief overview and the key characteristic of each family. The different families represent the possible ZLK dynamics. The initial conditions for each family and the parameter space are studies in detail in Grishin (2024a).
In Figure 2, we plot the correlation between $\omega$ and the GW frequency $f_{\rm GW}$ (Eq. 6) for the different families. Small libration occurs when the orbit is initialised with initial conditions at the fixed point (the small libration due to higher order effects, blue dots). For initial conditions further from the fixed point, but still ones that allow libration in $\omega$, the libration amplitude of $\omega$ and consequently of $e_{1}$ and $f_{\rm GW}$ is larger (black triangles). For an orbit outside the separatrix, $\omega$ completes a full circulation. The GW frequency becomes LISA-observable around $0.01$ Hz for some limited values of $\omega$ (red dots). For an initially librating orbit with sufficiently large eccentricity, GW dissipation at the pericentre significantly shifts the curves to a new energy curve. Eventually the orbit is tight enough and begins to decouple from the tertiary. At some point $\omega$ starts to circulate and eventually is dominated by GR 1PN precession without changing the eccentricity (green triangles). This orbit regularly achieves GW frequency above $0.1\ \rm Hz$.
@@
Figure 4: Representative of the large amplitude libration orbital family. Similar to Figure 3, but with different initial conditions (see table 1).
@@
Figure 6: Representative of the librating and merging family of orbits. The panels are similar to Figure 3. The initial conditions are described in table 1.
@@ 3.1 Circulating orbits
The only example of a population of highly inclined hierarchical triples with extreme mass ratios is the population of irregular satellites around giant planets in the Solar System. Out of the $228$ irregular satellites [^3] By irregular we mean satellites with orbit beyond the Laplace radius (Tremaine, 2023), which are believed to be formed by capture., only $13$ are found in ZLK libration, the rest are circulating (Grishin, 2024b). Theoretically, the area in $e_{1}$-$\omega$ phase space is larger for circulating orbit, with the actual ratio depending on $j_{z}$. Thus, the circulating family of orbits is the most common. Figure 3 depicts the orbital evolution of a representative orbit. The orbit starts from a relatively low/moderate eccentricity (in this case $e_{1}=0.5$, top left) and grows to large values. The mutual inclination is sharply decreasing in a correlated manner that preserves $j_{z}$ (on average; middle left). The argument of pericentre $\omega$ is circulating and completes a full revolution (bottom left).
The top middle panel shows the instantaneous separation ${\rm d}r$ and the pericentre $a_{1}(1-e_{1})$ which also changes by almost three orders of magnitude. The central panels show the evolution of the peak frequency $f_{\rm GW}$ with time. The peak frequency grows by almost three orders of magnitude for the circulating family. The middle bottom panel shows both strain polarisations $h_{+}$ and $h_{\times}$ for $D=8\ \rm kpc$, $\theta,\phi=0$ (head-on GW propagation). We see that the strain is highly bursty and peaks to strain magnitudes of about $10^{-19}$ for a short duration during the highly eccentric phase.
@@ 3.2 Large amplitude librating orbits
Figure 4 shows an example of an orbit that librates with a large amplitude $\omega$ between $\sim 0.75$ and $2.25$ rad, a range of almost $\pi/2$ radians (or 90 degrees) (top right). The only difference between this orbit and the circulating orbit is a different starting point of $\omega$. The peak frequency $f_{\rm GW}$ also changes by over two orders of magnitude, and the maximum peak is again achieved at $\omega\approx\pi/2$. The phase portrait of $e_{1}$-$\omega$ is now on a closed loop, but the behaviour is similar to the circulating orbit where $\omega$ sweeps quickly in the $e_{1}\sim\omega$ phase space. This is associated with the greatest loss of GW radiation, which is also indicated in the peak of the strain polarisations. Overall the large libration orbital family is similar to the circulating one. This is because these orbits are relatively close to the separatrix, which is the critical curve that separates circulating and librating orbits and crosses the unstable (hyperbolic) fixed point of $e_{1}=0,\omega=0$, similar to the pendulum. (see details e.g., in Grishin, 2024a).
@@ 3.3 Small amplitude librating orbits
Figure 5 shows the orbital evolution of a small amplitude librating orbit. We see that the orbital elements vary little, and the variation in $f_{\rm GW}$ varies by $0.4$ dex. The argument of pericentre is confined to a small range of $\sim 7.5$ degrees. The system stays at relatively high eccentricity, and the waveform has relatively regular bursts on the inner period timescale $P_{1}\approx 0.0092\ \rm yr$, which is different from the other systems that experience quiescent phases over the secular timescale of $\sim 1-2\ \rm yr$. The phase portrait shows clear motion of small librations around a fixed point. We will examine the small amplitude libration orbit, the different timescales involved (inner orbit $P_{1}\ll$ outer orbit $P_{2}\ll$ secular time $t_{\rm sec}$ and the comparison to an isolated binary evolution in sec. 4.4.